# FS-KAN: Permutation Equivariant Kolmogorov-Arnold Networks via Function Sharing

### v1

@@ 1 Introduction
From a theoretical perspective, we prove that FS-KAN architectures have equivalent expressive power to standard parameter-sharing MLPs in the uniform function approximation sense, implying that both architecture classes can represent the same functions. This equivalence establishes both fundamental limitations and guarantees for FS-KAN expressivity derived from equivalent results for parameter-sharing neural networks, including expressivity analysis for parameter-sharing-based GNNs (Chen et al., 2019; Geerts, 2020; Maron et al., 2019b, a; Azizian and Lelarge, 2020) and universality theorems for invariant or equivariant networks (Zaheer et al., 2017; Yarotsky, 2022; Segol and Lipman, 2019; Maron et al., 2019c; Keriven and Peyré, 2019) — all of which directly transfer using our result to their FS-KAN counterparts.
@@ 2 Related Work
Equivariance in deep learning. A function is equivariant to a group action if a group transformation of the input leads to a corresponding transformation of the output. Perhaps the most well-known example of this principle is convolutional neural networks (CNNs) (LeCun et al., 1998) that employ translation equivariant layers; more recent works have generalized equivariance to broader symmetry groups (Cohen and Welling, 2016; Ravanbakhsh et al., 2017; Qi et al., 2017; Zaheer et al., 2017; Esteves et al., 2018; Kondor et al., 2018; Weiler and Cesa, 2019; Pan and Kondor, 2022; Hands et al., 2024; Zhang et al., 2024). For instance, Zaheer et al. (2017) characterized permutation equivariant layers in the context of set inputs, paving the way for many set-based architectures such as Hartford et al. (2018); Maron et al. (2020). Kondor et al. (2018); Maron et al. (2019b, a); Pan and Kondor (2022) studied equivariant layers for graphs, and Cohen and Welling (2016) introduced a framework for building networks equivariant to additional transformations like rotations and reflections. Recent work by Navon et al. (2023); Zhou et al. (2023) leverages symmetries within neural networks to develop effective model editing and analysis techniques. As constraining models to be equivariant might lead to models with limited expressive power (Xu et al., 2019; Morris et al., 2019; Geerts, 2020), numerous works in recent years studied and characterized the expressive power of equivariant models (Zaheer et al., 2017; Maron et al., 2019a, c; Dym and Maron, 2020; Ravanbakhsh, 2020; Azizian and Lelarge, 2020).
@@ 3.2 Invariant and equivariant Function Sharing KA layers
Efficient FS-KA layers. In practice, standard KA layers can have significant computational and memory demands as they apply functions independently across all input and output pairs. However, in many cases, the sharing structure requires the same operations to be performed across multiple elements, followed by a sum aggregation. In linear layers, to reduce time and memory complexity, it can be helpful to commute sum or mean pooling with matrix multiplications. Inspired by this, we introduce the Efficient FS-KA layer, which first aggregates inputs according to the FS structure and then applies a shared KA sub-layer. While this relaxation does not yield equivalent layers, it reduces the number of nonlinear function applications while preserving equivariance, which can be easily verified. For example, consider a $S_{n}$-equivariant FS-KA layer (see equation 5). It applies a KA transformation independently to each element before performing sum-pooling over the set. In contrast, the efficient variant computes: $\tilde{\Phi}(\mathbf{x})_{q}=\tilde{\Phi}_{1}(\mathbf{x}_{q})+\tilde{\Phi}_{2}\left(\sum_{p=1}^{n}\mathbf{x}_{p}\right)$. In this variant, we sum over all elements so the second term can be computed once and reused (broadcast) for other computations within the layer. While the efficient FS-KA layer has the same number of parameters, it reduces memory usage, especially during training, by retaining smaller computational graphs. In general, the efficiency of the layer is determined by the group structure. We discuss on the efficiency and construction of efficient FS-KA layers for general permutation symmetry groups on Appx. A.3.
@@ 3.3 Examples and important cases
We now discuss specific instantiations of FS-KA layers to two key symmetry types: direct-product symmetries (Maron et al., 2020; Wang et al., 2020) and high-order tensor symmetries (Maron et al., 2019b, c; Kondor et al., 2018; Keriven and Peyré, 2019).
@@ 4 A theoretical analysis of the expressive power of FS-KANs
For any compact domain $\Omega$, $\epsilon>0$, and FS-KAN $\Phi$ with $l$ layers and continuous activations, there exists a parameter-sharing MLP $f$ with at most $2l$ ReLU layers s.t. $f-\Phi||_{\infty}<\epsilon$.
Both proofs provide constructive methods for building the corresponding networks, though not necessarily efficient in terms of depth or parameters. The results extend to higher-order tensors and direct products of symmetry groups via flattened tensor representations.
Implications for the expressivity of FS-KANs. These results establish that FS-KANs and parameter-sharing MLPs possess equivalent expressive power in the function approximation sense. This equivalence has profound implications for both theoretical analysis and practical applications. It allows us to transfer known properties and guarantees between these model classes, creating a bridge between the established theory of parameter-sharing networks and KANs. For example, if DeepSets is proven universal for permutation-invariant functions, we can establish that the corresponding FS-KAN construction inherits this universality. As a direct implication of the above-mentioned expressive power equivalence, we establish the following corollary,
@@ B.1 Key supporting lemmas
We begin by recalling Theorem 3.2 from Wang et al. [2024b] establishes that each layer of an MLP with activation function $\sigma_{k}(\cdot)=max(0,\cdot)^{k}$ can be represented by a KAN with two hidden layers and grid size $G=2$ with degree $k$ B-splines. Their proof demonstrates this by representing the linear part of an MLP layer using a single KAN layer and the non-linear activation using another KAN layer with a diagonal structure. i,e for a bounded domain $\Omega\subset\mathbb{R}^{d_{in}}$ and an affine layer $L:\mathbb{R}^{d_{in}}\to\mathbb{R}^{d_{out}}$ exists KA layers $\Phi:\mathbb{R}^{d_{in}}\to\mathbb{R}^{d_{out}}$ and $\Psi:\mathbb{R}^{d_{out}}\to\mathbb{R}^{d_{out}}$ such that:
@@
In the general case, the MLP has equivariant layers and invariant layer followed by a standard MLP. As the work of Wang et al. [2024b] already handles standard MLPs, we will focus on showing that any equivariant or invariant linear layer with $\sigma_{k}$ activation can be represented using one hidden layer FS-KAN.
Let $L:\Omega\to\mathbb{R}^{n\times d_{out}}$ be an equivariant affine layer. By Wang et al. [2024b] work exists a KA layer such that:
@@
Let $\epsilon>0$. By the universal approximation theorem Hornik et al. [1989] for MLPs with ReLU activations, for each function $\Phi^{(q,p)}$ (which is continuous on a compact domain), there exists a two-layer MLP $f^{(q,p)}$ with ReLU activations such that:
@@
$L_{1}$ is an equivariant linear map, so it must be a parameter-sharing linear layer [Wood and Shawe-Taylor, 1996]. Second layer construction $L_{2}:\mathbb{R}^{n^{2}\times d_{max}}\to\mathbb{R}^{n\times d_{out}}$ will be defined as:
@@
Using Lemma 6 from Lim et al. [2022] (Layer-wise universality implies universality), using composition of the MLPs approximation for each layer approximates the whole MLP up to any precision. ∎
@@ Appendix C Implementation details
All experiments on were implemented using the PyTorch framework and trained using the AdamW optimizer [Loshchilov and Hutter, 2017]. The FS-KAN variants implementations are based on the Efficient-KAN [Blealtan, 2024] package and the ConvKAN [Bodner et al., 2024] GitHub repository, with the exception of the symbolic formula experiments, which used the PyKAN library [Liu et al., 2024] and LBFGS optimizer for both KAN and FS-KAN implementations. For MLP-based models, we applied $\ell_{2}$ weight regularization; for KAN-based models, we used the regularization provided by the Efficient-KAN library. The exact parameter count of the invariant/equivariant models are in Table 2. As for the non-invariant standard KAN in the point cloud classification task, the parameter counts vary between 42,240 and 503,040 depending on the specific configuration. The regularization term was scaled by a hyper-parameter $\eta$ in both cases. Models were trained on NVIDIA V100 GPUs. We trained the models with 5 different seeds in all experiments and configurations.
While Group Convolutional Neural Networks (GCNNs) Cohen and Welling [2016] can also be considered as a baseline, they become computationally intractable for the large symmetry groups we consider, as group convolutions scale with the size of the group. This illustrates how our approach is more practical for real-world applications.
@@ C.1 Learning invariant symbolic formulas
The dataset generation follows the procedure provided in Liu et al. [2024]. FS-KAN models consist of one equivariant layer with single feature channels followed by one invariant layer. While the FS-KAN is illustrated for $n=3$, it can handle varying input sizes. For the baseline KAN, we use one hidden layer of width 3. Both models use L-BFGS optimization with 25 steps and $\lambda=0.001$. We refer to figure 7 for more results on different invariant formulas.
@@ C.2 Signal classification experiment
Data preparation. The dataset generation follows the procedure described on Maron et al. [2020], where each clean signal was randomly selected from a predefined set of types - sine, square, and sawtooth with $T=100$ time steps. To construct a set of noisy measurements, each signal is duplicated multiple times, adding independent Gaussian noise to each copy. All samples were generated using the code provided by Maron et al. [2020]. Validation and test sets have a fixed size of $300$ samples each.
Networks and training. FS-KAN and efficient FS-KAN models have 3 hidden layers of width $16,16,8$, while the baseline DSS used hidden layers of width $160,160,80$. Each layer was followed by a batch norm as well. We used the DSS implementation provided by Maron et al. [2020]. As mentioned earlier, we also trained a scaled DSS to match the number of parameters in the KAN-based models. We used a batch size of $64$ as in Maron et al. [2020]. For each configuration, we trained the models with 5 different seeds. We report in Table 3 the training durations for all models.
@@ C.3 Point clouds classification
Data preparation. For this experiment, we used ModelNet40 [Wu et al., 2015], which contains around $12\times 10^{3}$ point clouds from different $40$ classes. Each point cloud contains $n=1024$ points represented by $d=3$ spatial coordinates. We used the normalized and sampled point clouds by Qi et al. [2017] with no data augmentations. For the validation set, we used a $90\%-10\%$ split of the full train set.
Networks and training. All invariant models are composed of equivariant layers, followed by an invariant layer and a standard output layer of width $40$. FS-KAN and efficient KANs are composed of 2 equivariant FS-KA layers of width $36$, each followed by a $1D$ batch norm. The invariant layer has an output dimension of $36$ as well. The DeepSets model has the same architecture, only with a width of $128$ to match the number of parameters ($5.6\times 10^{4}$). The non-equivariant KAN consists of 3 hidden layers of width $16$, but since it cannot handle varying input sizes, its parameter count depends on the number of input points, ranging from $4.2\times 10^{4}$ to $5.0\times 10^{5}$ parameters. The transformer model uses an embedding dimension of $38$ with $3$ layers ($5.5\times 10^{4}$ parameters), implemented with an input projection layer, transformer encoder layers with multi-head attention and a classification head with intermediate dimensionality reduction. For training and testing, we used a batch size of 32. Additionally, we conducted an experiment to measure the average runtime and peak GPU memory usage for both the training and inference steps for the FS-KANs and DeepSets, as reported in Tables 4 and 5.
@@ C.5 Semi-supervised rating prediction experiment
Data preparation. We used the MovieLens 100k dataset [Harper and Konstan, 2015] along with smaller versions of Douban, Flixster, and Yahoo, as introduced by Monti et al. [2017]. In the Flixster dataset, where ratings are given in 0.5 increments, we labeled rating $0.5$ as $1$ represented non-integer ratings as probabilistic mixtures (e.g., a rating of 3.5 was modeled as a $50\%$ chance of 3 and $50\%$ chance of 4). For the Yahoo dataset, ratings ranging from 1 to 100 were clustered into 5 categories and treated as 151-5 values.
To simulate low-data scenarios, we uniformly sampled subsets of rows and columns from each rating matrix. Within each resulting sub-matrix, we further sub-sampled the observed entries to reduce training density. The train set was composed of uniformly sampled sub-matrices of the sparse matrix. The train set consists of 32 sub-matrices, and the test and validation sets have 64 samples each. The full data sampling process is illustrated in figure 8. Data was split into train, validation, and test sets using the U1 split defined by Harper and Konstan [2015].
Network and training. All models are composed of 9 equivariant layers, each followed by a skip connection and batch norm. KAN and FS-KAN architectures used layers of width 16 while SSEM uses layers of width 256 as detailed on Hartford et al. [2018]. A smaller SSEM with a width of 52 was also trained to match the parameter count of the KAN models. For each configuration, we trained the models with 5 different seeds.

### v2

@@ 1 Introduction
From a theoretical perspective, we prove that FS-KAN architectures have equivalent expressive power to standard parameter-sharing MLPs in the uniform function approximation sense, implying that both architecture classes can represent the same functions. This equivalence establishes both fundamental limitations and guarantees for FS-KAN expressivity derived from equivalent results for parameter-sharing neural networks, including expressivity analysis for parameter-sharing-based GNNs (Chen et al., 2019; Geerts, 2020; Maron et al., 2019b; a; Azizian and Lelarge, 2020) and universality theorems for invariant or equivariant networks (Zaheer et al., 2017; Yarotsky, 2022; Segol and Lipman, 2019; Maron et al., 2019c; Keriven and Peyré, 2019) — all of which directly transfer using our result to their FS-KAN counterparts.
@@ 2 Related Work
Equivariance in deep learning. A function is equivariant to a group action if a group transformation of the input leads to a corresponding transformation of the output. Perhaps the most well-known example of this principle is convolutional neural networks (CNNs) (LeCun et al., 1998) that employ translation equivariant layers; more recent works have generalized equivariance to broader symmetry groups (Cohen and Welling, 2016; Ravanbakhsh et al., 2017; Qi et al., 2017; Zaheer et al., 2017; Esteves et al., 2018; Kondor et al., 2018; Weiler and Cesa, 2019; Pan and Kondor, 2022; Hands et al., 2024; Zhang et al., 2024). For instance, Zaheer et al. (2017) characterized permutation equivariant layers in the context of set inputs, paving the way for many set-based architectures such as Hartford et al. (2018); Maron et al. (2020). Kondor et al. (2018); Maron et al. (2019b; a); Pan and Kondor (2022) studied equivariant layers for graphs, and Cohen and Welling (2016) introduced a framework for building networks equivariant to additional transformations like rotations and reflections. Recent work by Navon et al. (2023); Zhou et al. (2023) leverages symmetries within neural networks to develop effective model editing and analysis techniques. As constraining models to be equivariant might lead to models with limited expressive power (Xu et al., 2019; Morris et al., 2019; Geerts, 2020), numerous works in recent years studied and characterized the expressive power of equivariant models (Zaheer et al., 2017; Maron et al., 2019a; c; Dym and Maron, 2020; Ravanbakhsh, 2020; Azizian and Lelarge, 2020).
@@ 3.1 Preliminaries
Where the $\star$’ operator denotes applying the KA layer to the vector.
@@ 3.2 Invariant and equivariant Function Sharing KA layers
Efficient FS-KA layers. In practice, standard KA layers can have significant computational and memory demands as they apply functions independently across all input and output pairs. However, in many cases, the sharing structure requires the same operations to be performed across multiple elements, followed by a sum aggregation. In linear layers, to reduce time and memory complexity, it can be helpful to commute sum or mean pooling with matrix multiplications. Inspired by this, we introduce the Efficient FS-KA layer, which first aggregates inputs according to the FS structure and then applies a shared KA sub-layer. While this relaxation does not yield equivalent layers, it reduces the number of nonlinear function applications while preserving equivariance, which can be easily verified. For example, consider a $S_{n}$-equivariant FS-KA layer (see equation 5). It applies a KA transformation independently to each element before performing sum-pooling over the set. In contrast, the efficient variant computes: $\tilde{\Phi}(\mathbf{x})_{q}=\tilde{\Phi}_{1}(\mathbf{x}_{q})+\tilde{\Phi}_{2}\left(\sum_{p=1}^{n}\mathbf{x}_{p}\right)$. In this variant, we sum over all elements so the second term can be computed once and reused (broadcast) for other computations within the layer. While the efficient FS-KA layer has the same number of parameters, it reduces memory usage, especially during training, by retaining smaller computational graphs.
For arbitrary permutation groups $G$, efficient FS-KA layers are derived following the same principle: commuting element aggregations (sum or mean pooling) with application of shared functions according to the original sharing structure. In general, the efficiency of the layer is determined by the group structure. We discuss the efficiency and provide a more formal construction of efficient FS-KA layers for general permutation symmetry groups in Appx. A.3.
@@ 3.3 Examples and important cases
We now discuss specific instantiations of FS-KA layers to two key symmetry types: direct-product symmetries (Maron et al., 2020; Wang et al., 2020) and high-order tensor symmetries (Maron et al., 2019b; c; Kondor et al., 2018; Keriven and Peyré, 2019).
@@ 4 A theoretical analysis of the expressive power of FS-KANs
For any compact domain $\Omega$, $\epsilon>0$, and FS-KAN $\Phi$ with $l$ layers and continuous activations, there exists a parameter-sharing MLP $f_{\epsilon}$ with at most $2l$ ReLU layers s.t. $f-\Phi||_{\infty}<\epsilon$.
Both proofs provide methods for building the corresponding networks, though not necessarily efficient in terms of depth or parameters. The results extend to higher-order tensors and direct products of symmetry groups via flattened tensor representations.
Implications for the expressivity of FS-KANs. We show that for a specific permutation group $G$, FS-KANs and parameter-sharing MLPs can approximate each other arbitrarily well, meaning both model families have equivalent expressive power in the uniform approximation sense. This equivalence has profound implications for both theoretical analysis and practical applications. It allows us to transfer known properties and guarantees between these model classes, creating a bridge between the established theory of parameter-sharing networks and KANs. For example, if DeepSets is proven universal for permutation-invariant functions, we can establish that the corresponding FS-KAN construction inherits this universality. As a direct implication of the above-mentioned expressive power equivalence, we establish the following corollary,
@@ B.1 Key supporting lemmas
We begin by recalling Theorem 3.2 from Wang et al. (2024b) establishes that each layer of an MLP with activation function $\sigma_{k}(\cdot)=max(0,\cdot)^{k}$ can be represented by a KAN with two hidden layers and grid size $G=2$ with degree $k$ B-splines. Here, $k$ determines the polynomial degree of both the activation and the spline basis functions, while $G$ specifies the number of grid intervals used in the spline parameterization. Their proof demonstrates this by representing the linear part of an MLP layer using a single KAN layer and the non-linear activation using another KAN layer with a diagonal structure. i,e for a bounded domain $\Omega\subset\mathbb{R}^{d_{in}}$ and an affine layer $L:\mathbb{R}^{d_{in}}\to\mathbb{R}^{d_{out}}$ exists KA layers $\Phi:\mathbb{R}^{d_{in}}\to\mathbb{R}^{d_{out}}$ and $\Psi:\mathbb{R}^{d_{out}}\to\mathbb{R}^{d_{out}}$ such that:
@@
In the general case, the MLP has equivariant layers and invariant layer followed by a standard MLP. As the work of Wang et al. (2024b) already handles standard MLPs, we will focus on showing that any equivariant or invariant linear layer with $\sigma_{k}$ activation can be represented using one hidden layer FS-KAN.
Let $L:\Omega\to\mathbb{R}^{n\times d_{out}}$ be an equivariant affine layer. By Wang et al. (2024b) work exists a KA layer such that:
@@
Let $\epsilon>0$. By the universal approximation theorem Hornik et al. (1989) for MLPs with ReLU activations, for each function $\Phi^{(q,p)}$ (which is continuous on a compact domain), there exists a two-layer MLP $f^{(q,p)}$ with ReLU activations such that:
@@
$L_{1}$ is an equivariant linear map, so it must be a parameter-sharing linear layer (Wood and Shawe-Taylor, 1996). Second layer construction $L_{2}:\mathbb{R}^{n^{2}\times d_{max}}\to\mathbb{R}^{n\times d_{out}}$ will be defined as:
@@
Using Lemma 6 from Lim et al. (2022) (Layer-wise universality implies universality), using composition of the MLPs approximation for each layer approximates the whole MLP up to any precision. ∎
@@ Appendix C Implementation details
All experiments on were implemented using the PyTorch framework and trained using the AdamW optimizer (Loshchilov and Hutter, 2017). The FS-KAN variants implementations are based on the Efficient-KAN (Blealtan, 2024) package and the ConvKAN (Bodner et al., 2024) GitHub repository, with the exception of the symbolic formula experiments, which used the PyKAN library (Liu et al., 2024) and LBFGS optimizer for both KAN and FS-KAN implementations. For MLP-based models, we applied $\ell_{2}$ weight regularization; for KAN-based models, we used the regularization provided by the Efficient-KAN library. The exact parameter count of the invariant/equivariant models are in Table 2. As for the non-invariant standard KAN in the point cloud classification task, the parameter counts vary between 42,240 and 503,040 depending on the specific configuration. The regularization term was scaled by a hyper-parameter $\eta$ in both cases. Models were trained on NVIDIA V100 GPUs. We trained the models with 5 different seeds in all experiments and configurations.
While Group Convolutional Neural Networks (GCNNs) Cohen and Welling (2016) can also be considered as a baseline, they become computationally intractable for the large symmetry groups we consider, as group convolutions scale with the size of the group. This illustrates how our approach is more practical for real-world applications.
@@ C.1 Learning invariant symbolic formulas
The dataset generation follows the procedure provided in Liu et al. (2024). FS-KAN models consist of one equivariant layer with single feature channels followed by one invariant layer. While the FS-KAN is illustrated for $n=3$, it can handle varying input sizes. For the baseline KAN, we use one hidden layer of width 3. Both models use L-BFGS optimization with 25 steps and $\lambda=0.001$. We refer to figure 7 for more results on different invariant formulas.
@@ C.2 Signal classification experiment
Data preparation. The dataset generation follows the procedure described on Maron et al. (2020), where each clean signal was randomly selected from a predefined set of types - sine, square, and sawtooth with $T=100$ time steps. To construct a set of noisy measurements, each signal is duplicated multiple times, adding independent Gaussian noise to each copy. All samples were generated using the code provided by Maron et al. (2020). Validation and test sets have a fixed size of $300$ samples each.
Networks and training. FS-KAN and efficient FS-KAN models have 3 hidden layers of width $16,16,8$, while the baseline DSS used hidden layers of width $160,160,80$. Each layer was followed by a batch norm as well. We used the DSS implementation provided by Maron et al. (2020). As mentioned earlier, we also trained a scaled DSS to match the number of parameters in the KAN-based models. We used a batch size of $64$ as in Maron et al. (2020). For each configuration, we trained the models with 5 different seeds. We report in Table 3 the training durations for all models.
@@ C.3 Point clouds classification
Data preparation. For this experiment, we used ModelNet40 (Wu et al., 2015), which contains around $12\times 10^{3}$ point clouds from different $40$ classes. Each point cloud contains $n=1024$ points represented by $d=3$ spatial coordinates. We used the normalized and sampled point clouds by Qi et al. (2017) with no data augmentations. For the validation set, we used a $90\%-10\%$ split of the full train set.
Networks and training. All invariant models are composed of equivariant layers, followed by an invariant layer and a standard output layer of width $40$. FS-KAN and efficient KANs are composed of 2 equivariant FS-KA layers of width $36$, each followed by a $1D$ batch norm. The invariant layer has an output dimension of $36$ as well. The DeepSets model has the same architecture, only with a width of $128$ to match the number of parameters ($5.6\times 10^{4}$). The non-equivariant KAN consists of 3 hidden layers of width $16$, but since it cannot handle varying input sizes, its parameter count depends on the number of input points, ranging from $4.2\times 10^{4}$ to $5.0\times 10^{5}$ parameters. The transformer model uses an embedding dimension of $38$ with $3$ layers ($5.5\times 10^{4}$ parameters), implemented with an input projection layer, transformer encoder layers with multi-head attention and a classification head with intermediate dimensionality reduction. For training and testing, we used a batch size of 32. Additionally, we conducted an experiment to measure the average runtime and peak GPU memory usage for both the training and inference steps for the FS-KANs and DeepSets, as reported in Tables 4 and 5. We also conducted experiments in the high-data regime using the full ModelNet40 dataset (8,859 training samples with 1,024 points per cloud), a standard benchmark in 3D computer vision. The results, presented in Table 6, demonstrate that FS-KAN variants achieve comparable accuracy with established architectures like DeepSets and Point-Transformer in this setting, complementing our low-data regime analysis.
@@
Model DeepSets FS-KAN Efficient FS-KAN Point-Transformer Standard KAN
Test Accuracy $0.844\pm 0.005$ $0.837\pm 0.007$ $0.840\pm 0.005$ $0.840\pm 0.008$ $0.232\pm 0.074$
Table 6: Test accuracy on full ModelNet40 dataset (high-data regime)
@@ C.5 Semi-supervised rating prediction experiment
Data preparation. We used the MovieLens 100k dataset (Harper and Konstan, 2015) along with smaller versions of Douban, Flixster, and Yahoo, as introduced by Monti et al. (2017). In the Flixster dataset, where ratings are given in 0.5 increments, we labeled rating $0.5$ as $1$ represented non-integer ratings as probabilistic mixtures (e.g., a rating of 3.5 was modeled as a $50\%$ chance of 3 and $50\%$ chance of 4). For the Yahoo dataset, ratings ranging from 1 to 100 were clustered into 5 categories and treated as 151-5 values.
To simulate low-data scenarios, we uniformly sampled subsets of rows and columns from each rating matrix. Within each resulting sub-matrix, we further sub-sampled the observed entries to reduce training density. The train set was composed of uniformly sampled sub-matrices of the sparse matrix. The train set consists of 32 sub-matrices, and the test and validation sets have 64 samples each. The full data sampling process is illustrated in figure 8. Data was split into train, validation, and test sets using the U1 split defined by Harper and Konstan (2015).
Network and training. All models are composed of 9 equivariant layers, each followed by a skip connection and batch norm. KAN and FS-KAN architectures used layers of width 16 while SSEM uses layers of width 256 as detailed on Hartford et al. (2018). A smaller SSEM with a width of 52 was also trained to match the parameter count of the KAN models. For each configuration, we trained the models with 5 different seeds.